Shown below are data taken from a recent annual report of Reebok International Ltd.    (Dollar amounts in millions.)

                           January 1,     December 31,

                                                                  2017              2017

         Balance sheet data:

            Current assets...........................$ 714                $ 784

            Total assets.............................. 1,063            1,166

            Current liabilities.......................    256                   203

            Total liabilities.........................      372                   322

            Total stockholders' equity.........    691                   844

       

Income statement data:

            Net sales..............................................$1,822

            Gross profit..........................................     750

            Operating income                                     293

            Net income............................................. 175

         Based upon the above information, indicate the best answer in the space

         provided.

The current ratio at December 31, 2017 (rounded to the nearest tenth) is:

                        a.   2.4 to 1.                C.   3.9 to 1.

                        b.   .7 to 1.                  D. none

The gross profit rate for the year (rounded to the nearest 1 percent) was:

                        a.   41%.                     c.   64%.

                        b.   59%.                     d. none

The return on average total assets during the year (rounded to the nearest percent) was:

                        a.   16%.                     c.   67%.

                        b.   26%.                    d. none

The return on average total stockholders' equity during the year

                     (rounded to the nearest 1 percent) was:

                        a.   98%.                     c.   23%.

                        b.   38%.                     d. none
Answer #1

Current Ratio = Current Assets/Current Liabilities

=784/203

= 3.9:1

Hence, the answer is C.   3.9 to 1.

Gross Profit Rate = Gross Profit/Sales

= 750/1822

= 41%

Hence, the answer is a.   41%.

return on average total assets = Net Income/Average Total Assets

Average Total Assets = (1063+1166)/2 = 1,114.5

return on average total assets= 175/1114.5

= 15.7% or 16%

Hence, the answer is      a.   16%.

return on average total stockholders' equity = Net Income/Average Equity

Average Equity = (691+844)/2 = 767.5

return on average total stockholders' equity = 175/767.5

= 22.8% or 23%

Hence, the answer is   c.   23%.